OpenStudy (anonymous):

2. An object is initially at rest and starts to accelerate uniformly at 3 m/s2 at time t = 0 s. At time t = 8 s, what would be the distance travelled by the object?

OpenStudy (anonymous):

we havethe formula: v=u+at,but u=0 so v=at ie v=24m/s again we have v^2=u^2+2as also u=0 so v^2=2as so s=v^2\[\div\]2a so we get s=96m

OpenStudy (anonymous):

s=ut+.5 at^2 u= o m/s a=3 m/s^2 t=8 s hence s=.5*3*8^2=96m